OpenStudy (anonymous):

evaluate limit as x goes to 2 = sin(x-2)/(x^2-4)

OpenStudy (anonymous):

infinity

OpenStudy (anonymous):

take the limit of top and bottom sin(0) = 1 (2^2 - 4) aproaches 0 BUT can never go to zero because you cant divide by zero

OpenStudy (anonymous):

so as the denominator goes to 0.00000001 or 0.000000000000000001

OpenStudy (anonymous):

its ilike this 1/4 then 1/2 then 1/1 then 1/0.000001 then 1/0.0000000000001

OpenStudy (anonymous):

This one can be handled with a little manipulation. The bottom can be expanded to (x-2)(x+2). The lim of [sin (x-2)]/(x-2)]=1. That leaves 1/(x+2). So lim is 1/4.

OpenStudy (anonymous):

no the bottom cannot be expanded like that

OpenStudy (anonymous):

I am guessing by the way these questions are set up that x^2-4 is not sin of, but just a number.

OpenStudy (anonymous):

I think the its sin of x-2 so sin(x-2)

OpenStudy (anonymous):

divided by x^2 - 4

OpenStudy (anonymous):

In that case we are saying the same thing. \[x ^{2}-4=(x-2)(x+2)\]The difference of two squares.

OpenStudy (anonymous):

no thats what you are saying

OpenStudy (anonymous):

that is impossible to do

OpenStudy (anonymous):

I have had a couple beers. But are you telling me that the above is not the difference of two squares?

OpenStudy (anonymous):

no it is

OpenStudy (anonymous):

maybe your right im sorry

OpenStudy (anonymous):

limit as x goes to 2 = sin(x-2)/(x^2-4) 0/0 so Use L'Hopital's rule lim x->2 Cos(x-2)/(2x) = Cos (0) / (2*2) = 1/4